zakary429:

Makayla is taking a math course and is working with the perimeter of rectangles. She knows the perimeter and length of her rectangle but wants to solve for the width. Rearrange the following equation for W, where P is the perimeter, L is the length, and W is the width of the rectangle. P = 2L + 2W W = 2P − 2L W equals P over 2 minus 2 times L W equals the quantity P minus 2 times L all over 2 W = 2P + 2L

zakary429:

its part of the equation

zakary429:

could you help me solve the equation

simsharrison:

Yes, to make the equation easier plug in random numbers to see what you get and then plug in numbers into the answers choices *same number, same spot* it will take longer but you will get the answer

zakary429:

i will try that thanks

simsharrison:

@Convert I know the answer, I just dont know how to put it into words, like walk them through

Convert:

Ok so P = 2L + 2W is the formula for perimeter So using this we have to solve for W so we use a simple algebraic expression: P = 2L + 2W -2W -2W -2W + P = 2L -P -P -2W = -P + 2L -2W/-2 = (-P/2) + (2L/-2) W = -P - 2L all dived by 2 Answer: 3rd option

Convert:

Sorry I had to make the spacing different so the equal sign would allign

simsharrison:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert Ok so P = 2L + 2W is the formula for perimeter So using this we have to solve for W so we use a simple algebraic expression: P = 2L + 2W -2W -2W -2W + P = 2L -P -P -2W = -P + 2L -2W/-2 = (-P/2) + (2L/-2) W = -P - 2L all dived by 2 Answer: 3rd option \(\color{#0cbb34}{\text{End of Quote}}\) I knew it was C, thanks I just didn't know how to put it into words

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert Ok so P = 2L + 2W is the formula for perimeter So using this we have to solve for W so we use a simple algebraic expression: P = 2L + 2W -2W -2W -2W + P = 2L -P -P -2W = -P + 2L -2W/-2 = (-P/2) + (2L/-2) W = -P - 2L all dived by 2 Answer: 3rd option \(\color{#0cbb34}{\text{End of Quote}}\) I knew it was C, thanks I just didn't know how to put it into words \(\color{#0cbb34}{\text{End of Quote}}\) Oh lol well there you go 😂

Convert:

And I meant divided* not dived

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert Ok so P = 2L + 2W is the formula for perimeter So using this we have to solve for W so we use a simple algebraic expression: P = 2L + 2W -2W -2W -2W + P = 2L -P -P -2W = -P + 2L -2W/-2 = (-P/2) + (2L/-2) W = -P - 2L all dived by 2 Answer: 3rd option \(\color{#0cbb34}{\text{End of Quote}}\) I knew it was C, thanks I just didn't know how to put it into words \(\color{#0cbb34}{\text{End of Quote}}\) Oh lol well there you go 😂 \(\color{#0cbb34}{\text{End of Quote}}\) ATTENTION please -P/(-2) not equal -P you missed from (-P/2) one minus sign of 2 please correct it ...

jhonyy9:

and -2W = -P + 2L -2W/-2 = (-P/2) + (2L/-2) W = -P - 2L all dived by 2 here not divided by 2 bc. you divide them by -2, minus 2 @Convert Attention please !

Seviackerman:

Please I need some help, here's the worded problem: Find the dimensions of the largest rectangle that can inscribed in the right triangle with sides 3, 4 and 5 (a) two sides of the rectangle are on the legs of the triangle, and if (b) a side of the rectangle is on the hypotenuse of the triangle. Thankyou in advance!!

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 and -2W = -P + 2L -2W/-2 = (-P/2) + (2L/-2) W = -P - 2L all dived by 2 here not divided by 2 bc. you divide them by -2, minus 2 @Convert Attention please ! \(\color{#0cbb34}{\text{End of Quote}}\) Oh thank you @jhonyy9 for pointing that out! I completely missed it at first my bad-